OpenStudy (danieke):

The price of a notebook has risen to $3.45 today. Yesterday's price was $3.20. Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (mathstudent55):

Remember the steps. 1. what is the change in price? subtract the old price from the new price

OpenStudy (rocketisawesome):

What class are you taking?. Because it is weird I have done these questions.

OpenStudy (danieke):

college readiness

OpenStudy (mathstudent55):

2. Now find what percent of the old price is the change divide the change by the old price and multiply by 100

OpenStudy (rocketisawesome):

the answer is

OpenStudy (rocketisawesome):

7.2% percent decrease

OpenStudy (danieke):

so its 3.20-3.45

OpenStudy (mathstudent55):

No. Subtract the old price FROM the new price. That means new price - old price 3.45 - 3.20

OpenStudy (mathstudent55):

$3.45 is today's price. This is the new price. $3.20 is yesterday's price. This is the old price. The change in price is the new price minus the old price. Change in price = 3.45 - 3.20

OpenStudy (mathstudent55):

What do you get for the change in price?

OpenStudy (danieke):

0.25

OpenStudy (mathstudent55):

Correct. Now we do step 2 above. We need to find what percent of 3.20 is 0.25

OpenStudy (rocketisawesome):

Well, it looks like you are in good hands bye. I would help more but it looks like you have someone else helping you better then me

OpenStudy (mathstudent55):

To find the percent, we divide the change by the old price and multiply by 100. What is \(\dfrac{0.25}{3.20} \times 100 \) ?

OpenStudy (rocketisawesome):

Bye

OpenStudy (danieke):

7.8125

OpenStudy (mathstudent55):

Correct. Now round to nearest tenth of a percent.

OpenStudy (danieke):

7.8 thank uu so much guys

OpenStudy (mathstudent55):

7.8% is correct

OpenStudy (mathstudent55):

You're welcome.